OpenStudy (anonymous):

Help 8 – 3x – x = 12 x =

OpenStudy (serenity74):

Okay m8 so you have to isolate the numbers from the equation and solve for x.

OpenStudy (serenity74):

so what do you do?

OpenStudy (anonymous):

add 8 on both sides?

OpenStudy (serenity74):

no subtract since 8 is positive

OpenStudy (anonymous):

oh

OpenStudy (serenity74):

yes then what do you do?

OpenStudy (anonymous):

I got 4 so the equation looks like this now? 3x-x=4

OpenStudy (serenity74):

No lol you forgot the -3x -x = 4 -4x = 4 x = -1 8 was positive all along but 3x was negative since the minus sign was in front of it.

OpenStudy (anonymous):

3x + -x?

OpenStudy (serenity74):

Do you get it? If not let me know and if you have anymore questions just tag me and I will try my best to help you.

OpenStudy (serenity74):

wait what? the question said 8 - 3x -x = 12 so you isolate the x which means you subtract 8 on both sides which leaves you with -3x - x = 12 - 8 then you combine like terms -4x = 4 then solve for x x = 4/-4 x = -1

OpenStudy (anonymous):

how did you get the negative sign?

OpenStudy (serenity74):

at which part? I just copied what your question said

OpenStudy (anonymous):

-3x - x = 12 - 8

OpenStudy (serenity74):

ohhh you mean how did I get -4x? -3x -x = -4x because when you add two minus signs together you get a negative answer -8 because you subtracting it from the 12 side

OpenStudy (serenity74):

you subtracted it on both sides right? 8 = 12 -8 -8 _____ ______ 0 4

OpenStudy (anonymous):

yeah, which comes out to 3x-x=4?

OpenStudy (serenity74):

no you still have -3x -x = 4 -4x => since negative + negative = negative answer

OpenStudy (serenity74):

-3x + -x = -3x -x = -4x

OpenStudy (anonymous):

I know but how did you get the negative sign? Once subtracting from both sides minus 8 cancels out

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

or am I missing something

OpenStudy (serenity74):

Yes on the left side not on the right side

OpenStudy (serenity74):

Whatever you do on the left you have to do the same thing on the right hence why you are subtracting 8 from 12 on the right side

OpenStudy (anonymous):

right, so that comes out as 4

OpenStudy (serenity74):

yes

OpenStudy (serenity74):

but you still have to solve for X and you still have -3x - x = 4 so you have to do next is combine like terms => -4x = 4 then you divide the -4 on both sides which gives you x = -1

OpenStudy (anonymous):

wait so if I subtracted -8 from both sides like this, 3x -8-x=12 -8 -8 it'd come out as positive?

OpenStudy (anonymous):

but if I did it on the right side 8 - 3x - x =12 it would be negative?

OpenStudy (serenity74):

YES lol

OpenStudy (anonymous):

Ohhh

OpenStudy (serenity74):

I am so confused wait what's the question?

OpenStudy (serenity74):

The real one so i can see whether what you are talking cause that matters a lot

OpenStudy (anonymous):

???? \[8-3x-x=12\]

OpenStudy (anonymous):

it's been 20 minutes ;-;

OpenStudy (serenity74):

okay that's the question right?

OpenStudy (anonymous):

yes.

OpenStudy (serenity74):

Dude you keep messing me up and the answer is a negative answer

OpenStudy (anonymous):

I was just wondering why it was a negative number instead of positive -_-

OpenStudy (serenity74):

Well that's because you are dividing it by a negative number

OpenStudy (anonymous):

dividing? where did that come from

OpenStudy (serenity74):

well sorry I tried to explain it clearly m8 don't get all grumpy at me :p sorry if i came off rude

OpenStudy (serenity74):

-4x = 4 x = 4 / (-4) = - 1

OpenStudy (anonymous):

ok well thanks for the help

OpenStudy (serenity74):

sorry if i confused you dude sorry

OpenStudy (anonymous):

it's fine, I didn't think you were rude at all

OpenStudy (serenity74):

:)